OpenStudy (anonymous):

Which is equivalent to (-2 + 4i)(7 + 2i)

OpenStudy (anonymous):

-14-4i+28i+8i^2 -14+24i-8 -22+24i

OpenStudy (callisto):

i guess i saw the same question before?

OpenStudy (tommo_lcfc):

Expanding the brackets gives -14 + 28i - 4i + 8i^2 = -14 + 24i + 8(-1) = -14 + 24i -8 = 24i - 22 or -22 + 24i